The radius of a circle is 18. Find the length of the peri

meter and the apothem of a

regular

(a) inscribed quadrilateral,

and

(b) circumscribed hexagon.



1
Expert's answer
2021-01-29T12:11:24-0500

An apothem of a regular polygon will always be a radius of the inscribed circle. It is also the minimum distance between any side of the polygon and its center.



(a) Consider isosceles right triangle "N_1OA"


"OA=AN_1=\dfrac{R}{\sqrt{2}}=\dfrac{18}{\sqrt{2}}=9\sqrt{2}"

"N_1N_4=2AN_1=18\sqrt{2}"

"Perimeter=4N_1N_4=72\sqrt{2}"

"apothem=OA=9\sqrt{2}"

(b)  The circumscribed hexagon is defined as the hexagon is outside in the circle or circle is inside in the hexagon.

Consider  equilateral triangle "M_1OM_6"


"\angle M_1OM_6=60\degree"

Consider right triangle "M_1OB"


"OB=R=18, \angle M_1OB=30\degree"

"OM_1=\dfrac{OB}{\cos\angle M_1OB}=\dfrac{18}{\sqrt{3}/2}=12\sqrt{3}"

"M_1M_6=OM_1=12\sqrt{3}"

"Perimeter=6M_1M_6=72\sqrt{3}"

"apothem=OB=18"